Hobart recoded to 2 types of disk media. Xraid @ 256 Mbps AND Mark5a @ 512 Mbps.
Xraid data on curtV012A1-7_p2 first 12 hrs then curtV012A1-7_p3 for second 12 hours. This disk set also has data from experiment v252h and so will be sent to Curtin.
Mark5A diskpack MPI-0519/4000/1024 will be sent to Bonn using DHL once Chris says. For the moment the Mark5A disk pack will stay in the recorder at Hobart. This disk pack also contains the VC067 fringe check experiment data at the start. 3.3TB total were recorded on the Mark5 module. 3,283,765,272,448 bytes. This disk pack also contains the VC067 fringe check experiment data at the start.
Between the times of 223/2238 till 223/2335 the serial comms to the receiver were hung up and there was no tsys data AND the noise diode was ON.
Observers: Brett, Jamie M, Jay Blanchard
2008.222.23:59:42.03;“Injecting test tones to ckck the mk5 set up.
2008.223.00:05:50.71;”System coherent, tones off
2008.223.00:05:56.18;“System coherent, tones off.
2008.223.00:06:17.68;”All VC frequencies are correct
2008.223.00:14:09.52;“Weather is cold, calm and overcast.
2008.223.04:54:06.63;“VC levels fine
2008.223.04:54:59.32;”Weather is cold and overcast, with light winds.
2008.223.05:36:26.50;“first 12h of LBA data recorded CURT_V012A_P2
2008.223.05:36:49.41;”second 12h of LBA data will be recorded yo CURT_V012A_P3
2008.223.12:01:14.77;“disk change to l_3 (CURT_V012A_3)
2008.223.23:31:01.73;”noise diodes have been on. Restarting Rakbus
2008.223.23:37:07.06;“just cycled power to S/X receiver. Noise diode is now under control again
2008.223.23:57:25.73;”for the hour before recycling the receiver power, the Tsys measurements and cryo measurement are rubbish and the noise cal diode was stuck ON!
The following lines were the edited lines to change the frequencies so there was no repatching:
Following are the original lines from v271a.prc that were edited above: